

********************************************************************
UARS WEEKLY STATUS REPORT - 01 MAR 2002
********************************************************************
***********
SOLSTICE
***********
Nothing received.
***********
SUSIM
***********
SUSIM performed its daily mid-resolution and MgII solar scans
during periods of predicted UARS telemetry.
The SUSIM data has been processed on the SUSIM RAC hosts for all
of the backlogged days from 1 Nov. 2001 to the present to level
3AS.
We are preparing nonfinal calibration determinations that will be
applied to recent UV irradiance data for use in comparisons with
the measurements of the TIMED/SEE and its calibration rocket.
On February 20, a test microload was sent from the ECLIPSE system
to SUSIM aboard UARS. The microload was successfully received by
the instrument with no problems.
We are evaluating the latest version of the CECIL-compatible
procedures derived in an automated fashion from the SUSIM STOL
procedures. Because SUSIM commanding depends more upon microloads
and less on STOL procedures and commands from the OBC, we don't
expect serious problems in the transition to the ECLIPSE system.
***********
ACRIM
***********
Nothing received.
***********
PEM
***********
A detailed review of the PEM CECIL PROCs were conducted. 10 of 51 PROCs
were found to be missing. GSFC has found these routines and sent them on to
PEM for review.
PEM continues to receive NMC data files from NOAA. No problems have
been encountered.
PEM awaits support of the UARS Traceability Mission for the TIMED mission.
Initial meetings with a GUVI team representative have highlighted the need
for knowledge of conjunction periods. Investigation is underway into
identifying times when the PEM spectrometers and the GUVI view plane are
sampling from the same volume of the atmosphere. To date, orbital analysis
has shown that during March, orbital intersection occurs at low-latitudes,
which are not useful time periods. Orbital investigations are now concentrating
on the month of April.
*Beginning of power orbit day (PEM activation): array current >5 amps, array
voltage >53.5 volts, and battery current >2.0 amps.
End of power orbit day (PEM deactivation): time in power orbit day >4 minutes,
array current <9 amps, and battery currents <0.2 amps.
Power orbit day variation: ~39-40 min at yaw to ~50-51 min at maximum beta.
***********
HALOE
***********
Nothing received.
***********
MLS
***********
Nothing received.
***********
HRDI
***********
Nothing received.
***********
WINDII
***********
Nothing received.
***********
MPG
***********
The MPG participated in the UARS ECLIPSE System
Meeting on 28 February 2002.
The MPG continues to serve as the point-of-contact
for UARS instrument operations to support re-engineering
of the UARS Mission Operations Concept to develop
capabilities for the Science Traceability Mission.
Daily processing of instrument Daily Activity Plans
continues with no problems.
***********
CDHF
***********
Nothing received.
***********
FOT
***********
REPORTING PERIOD
This report covers 16 February 2002 (Orbit 57048, GMT Day 047) through
22 February 2002 (Orbit 57152, GMT Day 053).
SPACECRAFT OPERATION
The observatory is now in Forward Flight (Southern Hemisphere viewing)
and is performing nominally under two-battery operations. The
instrument operational changes were:
INSTRUMENT OPERATIONAL CHANGES
Instrument Time Orbit Comment
-None this report period.
The beta angle ranged between the angles of 42.87 degrees to 44.43
degrees for this report period. The beta angle is now decreasing
towards a minimum angle of 0 degrees on 08 March 2002 (DOY 067).
Spacecraft battery 1, 2, and 3 performance monitoring continues.
Battery 1 remains OFF the charge relay. Battery 2 and Battery 3 are
currently maintaining greater than 26.24 V end-of-night (EON) load bus
voltage. Battery temperatures are stable with a temperature delta
between Battery 2 and Battery 3 of 1.31 to 1.31 degrees C. Additional
battery performance data for each battery on 16 February 2002 and 22
February 2002 are provided in Appendix D.
The clock error ranged from -11.1 to 10.8 msec during this report
period. Clock rate adjustments are listed in Appendix C.
All instrument and subsystem engineers were advised on the initial
excursion of ALERTS or Out-Of-Limits conditions detected by the FOT
for this report. A summary is listed below and the details of ALERT
and Out-Of-Limits occurrences are listed in Appendix B.
OUT-OF-LIMITS = 00
ALERTS = 00
The following real-time command sheets were executed for the listed
operational element(s) during this report period (see Appendix A).
FOT = 2
UARS Anomaly Reports initiated during this report period are listed in
Appendix E and summarized below.
FOT = 0
Two (2) Anomaly Reports remained OPEN at the end of this report period
(see Appendix E).
TRANSPONDER FREQUENCIES:
Transponder A: 01/333 - 2287.496855
Transponder B: 02/002 - 2287.498845
GROUND SYSTEM OPERATION
Zero (0) operational support problems occurred this period. Anomaly
Reports are included in Appendix E.
The data loss calculations are no longer being supplied by Data
Capture Facility (DCF).
OTHER SIGNIFICANT EVENTS
UARS SATELLITE OPERATIONS
The second run of the day in the life (DITL) of UARS took place this
week. This test was setup using the UTTS and Eclipse to simulate
routine operations. The online operators sent loads to the simulator
just as they would to the spacecraft in order to find problems with
the Eclipse system and in order to get more familiar with Eclipse
operations.
On 19 February 2002 a test microprocessor load was sent from Eclipse
to the spacecraft. This microprocessor load contained three commands
for the SUSIM instrument. The command load was a success and the FOT
would like to thank the SUSIM team for volunteering for the test.
MEETINGS
The FOT attended a UARS Eclipse Working Meeting on Thursday, February
21, 2002.
FUTURE SIGNIFICANT EVENTS
The next UARS Yaw slew is scheduled for 08 March 2002 (DOY 067). This
maneuver will be a Forward to Reverse Flight maneuver (Northern
Hemisphere viewing).
Attachments: Appendix A, Real-time Command Sheets
Appendix B, Out-Of-Limits & ALERT Occurrences
Appendix C, Clock & Solar Array Rate Corrections
Appendix D, UARS Battery Performance Data
Appendix E, UARS Anomaly Reports
APPENDIX A
REAL-TIME COMMAND SHEETS
Week ending 22 February 2002
ORBIT AOS NO. SUBSYS REASON
----- --- --- ------ ------
57049 047/0221 02-058 OBC Clk Adj fr 20.6 to 0.0 msec/day
57095 050/0331 02-059 OBC Clk Adj fr 0.0 to 20.6 msec/day
APPENDIX B
OUT-OF-LIMITS & ALERTS OCCURRENCES
Week ending 22 February 2002
The following Out-Of-Limits occurrences were detected:
MNEMONIC STATE DESCRIPTION ORBITS
-------- ----- ----------- ------
-None this report period.
The following ALERT occurrences were detected:
MNEMONIC DESCRIPTION ORBITS
-------- ----------- ------
-None this report period.
_____________________________________________________________________
APPENDIX C
CLOCK & SOLAR ARRAY RATE CORRECTIONS
Week ending 22 February 2002
CLOCK DRIFT RATE ADJUSTMENTS
ORBIT TIME CLOCK RATE CHANGE
----- ----- -----------------
57049 047/0221 Clk Adj fr 20.6 to 0.0 msec/day
57095 050/0331 Clk Adj fr 0.0 to 20.6 msec/day
APPENDIX D
UARS BATTERY PERFORMANCE DATA
Week ending 22 February 2002
16 February 2002 - GMT Day 047 Beta = 42.87 deg, SA Pos = 269 deg
DIFF V (mV) TEMP (C) CURR (amp) EON V AVE C/D AVE
BAT MAX/MIN MAX/MIN MAX/MIN MIN RATIO DOD(%)
--- ------- -------- ---------- ----- ------- ------
1 -44.8/-56.0 -3.16/-4.38 +0.0/ -0.0 10.72 0.000 0.00
2 +11.2/-22.4 +5.14/+4.15 +26.4/-14.4 26.40 1.040 19.57
3 +16.8/-22.4 +3.82/+3.17 +25.6/-14.4 26.24 1.030 19.22
22 February 2002 - GMT Day 053 Beta = 44.43 deg, SA Pos = 269 deg
DIFF V (mV) TEMP (C) CURR (amp) EON V AVE C/D AVE
BAT MAX/MIN MAX/MIN MAX/MIN MIN RATIO DOD(%)
--- ------- ---------- ---------- ----- ------- ------
1 -50.4/-67.2 -3.16/-4.38 +0.0/ -0.0 10.72 0.000 0.00
2 +11.2/-22.4 +5.14/+4.15 +25.6/-14.4 26.24 1.040 19.69
3 +22.4/-22.4 +3.82/+3.17 +25.2/-14.4 26.24 1.030 19.34
V/T LEVEL CHANGES
ORB# TIME LVL FR-TO
----- ---- ---------
-no changes V/T 5
APPENDIX E
UARS ANOMALY REPORTS
Week ending 22 February 2002
NEW ANOMALY REPORTS GENERATED
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
-None this report period.
ANOMALY CLOSURE INFORMATION RECEIVED
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
-None this report period.
ANOMALY INVESTIGATION REPORTS GENERATED
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
-None this report period.
OPEN ANOMALY REPORTS
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
02-004 SOLSTICE 56883 036/0050 N/A Solstice in Wrong Mode
02-005 FOT see list Feb 02 14170 Generic Late Acquisition